I need to find the equation of a line through (-5,4) and parallel to -6x+4y=-7
The equation needs to be written in standard form.
I have tried to work several versions of this problem and get them wrong each time. I have problems half way through it, due to the fractions. My book is not clear enough.
Thank you

Answer by mananth(16949)

I need to find the equation of a line through (-5,4) and parallel to -6x+4y=-7
..
If books were very clear then what will teachers do?!!!!!
Anyway
...
the equation -6x+4y=-7
1.write in slope intercept form
4y=6x-7
2.compare with equation y = mx+b
slope m= 6
3. point (-5,4) m=6
4. find value of b by plugging x,y,m
4=6*-5+b
4=-30+b
b=34
5. the equation you need is parallel to the given line.
so slope is same. m=6
6. you have slope,m you have b, plug the values &
write the equation
y=6x+34
you want to write in general form. bring all terms on one side of the equation.
bring y to the RHS.
6x-y+4=0 is the required equation.
...
m.ananth@hotmail.ca